OpenStudy (anonymous):

whats the lim x goes to infinity of 2/square root of x

OpenStudy (anonymous):

0

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Rose if you're not convinced, just plug a large number into your calculator to check! :) if you plug one bajillion in for x, what value do you get for the limit? \[\frac{ 2 }{ \sqrt{bajillion} }\approx \frac{ 2 }{ 999999999 } \approx 0\]

OpenStudy (anonymous):

yeah, but do what zepdrix said only for check, because if you do this in a test for example, it will be wrong because infinity is NOT a number...

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

it is always 0 ?

OpenStudy (anonymous):

no! it will depends on what you are taking the limit, always try to simplify the limit, try to do not have a denominator, in this case we can not simplify so we do this

OpenStudy (anonymous):

i am trying to solve this problem where x goes to infinity 2+square root of x/ 2- square root of x

OpenStudy (anonymous):

i am getting the answer as 1, is it right?

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}\frac{2+\sqrt{x}}{2-\sqrt{x}}\]

OpenStudy (anonymous):

this one?

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

the limit is equal to -1

OpenStudy (anonymous):

have you already study L'hospital rule ?

OpenStudy (anonymous):

can you please explain how you got -1?

OpenStudy (anonymous):

No,

OpenStudy (anonymous):

ok let me explain

OpenStudy (anonymous):